OpenStudy (anonymous):

Derive the trigonometric addition formula for sine: sin(a + b) = sin a cos b + cos a sin b

OpenStudy (rogue):

Instead of expanding the thing out, you can also just apply the chain rule.\[\frac {d}{dx} \left[ \sin(a + b) \right] = \cos(a + b)\times \left[ \frac {da}{dx} + \frac {db}{dx} \right]\]

OpenStudy (anonymous):

i think the question asks to come up with this identity... nothing to do with derivatives..

OpenStudy (rogue):

oh, lol. I misread :(

OpenStudy (rogue):

Try this: http://www-spof.gsfc.nasa.gov/stargaze/Strig5.htm

OpenStudy (anonymous):

wow, nasa...

OpenStudy (rogue):

NASA scientists do lots of stuff when they're bored... http://apod.nasa.gov/htmltest/gifcity/e.1mil But the information databases they create are very useful =)